# Prediction of the Most Fire-Sensitive Point in Building Structures with Differentiable Agents for Thermal Simulators

### v1

@@ IV-G Summary of the Dataset Generation
Figure 5: Histograms of MIDR for labeled and unlabeled structures with gravity loads and fire cases.
V Maximum Interstory Drift Ratio Predictor
The MIDR Predictor functions as a differentiable agent for the FEA simulators. It accepts building structure data and fire coordinates as inputs and predicts the MIDR of the building under fire exposure. Leveraging GNNs for structural representation, the predictor integrates Transfer Learning (TL) to optimize data utilization while focusing on MIDR prediction. Additionally, an innovative Edge Update (EU) module, to update the edge attributes of the GNNs, is introduced to capture changes in element attributes during fire scenarios.
V-A Graph Neural Networks
V-A1 Overview of GNNs
Building structures and graphs share an intrinsic similarity: structural nodes correspond to graph nodes, and structural elements (e.g., beams and columns) map naturally to graph edges. This analogy forms the basis for utilizing GNNs in structural data processing.
A typical GNN consists of multiple stacked layers, where the output of one layer serves as the input for the next. Each layer executes three core operations –Message Passing, Aggregation, and Update– which can be expressed as follows:
where $\gamma^{k}(\cdot)$ and $\phi^{k}(\cdot)$ represent the update and message functions, respectively, $\bigoplus$ is the aggregation operation, e.g., sum, mean, or max, $\mathcal{N}(i)$ denotes the set of neighboring nodes of node $i$, and $\boldsymbol{v}_{i}^{k}$ and $\boldsymbol{e}_{j,i}$ are the attributes of node $i$ at the $k$-th layer and of the edge from node $j$ to node $i$, respectively. Here, both $\gamma^{k}(\cdot)$ and $\phi^{k}(\cdot)$ are typically chosen as Multi-Layer Perceptrons (MLPs), i.e., fully connected neural networks.
Figure 6 provides a schematic of a GNN with $K$ layers, illustrating the role of node and edge encoders, as well as the EU mechanism. The shown GNN incorporates initial node and edge encoders to transform input features into higher-dimensional spaces and utilizes an EU mechanism to dynamically update edge attributes during each layer to capture the changes of the elements in the fire scenario.
Figure 6: Illustration of a $K$-layer GNN architecture. The figure demonstrates information flow within the network for a graph with $N$ nodes.
V-A2 Handling structural data
A significant challenge in GNNs is the over-smoothing, where during each layer’s update process, a node’s attributes aggregate information from its neighboring nodes. Consequently, increasing the number of layers causes node attributes to homogenize, losing their specificity and negatively impacting the model’s performance. This issue is particularly pronounced in building structures with varying story counts and number of nodes, such as 2-story vs. 7-story buildings. If both are passed through a 7-layer GNN, the node attributes of the 2-story building would experience severe over-smoothing. To mitigate this problem, we adopt the method from [5], restricting the number of GNN layers for each graph, i.e., building structure herein, to the building’s maximum number of stories, ensuring critical information about fire propagation is retained in the node attributes without over-smoothing.
Beams and columns exhibit distinct load characteristics that influence how they are represented in graph-based model of a structure. Beams primarily bear gravity loads that act perpendicular to their axial direction, which aligns with their structural function in resisting bending moments. Under thermal expansion due to fire, these loads affect both end nodes of the beam. As a result, when representing beams in a graph model, it is intuitive to treat them as undirected edges, reflecting their bidirectional influence on the connected nodes. Columns, on the other hand, experience gravity loads along their axial direction, typically directed downward, which is consistent with their role in supporting vertical loads. When a column deforms under loads (due to thermal expansion, lateral loads, or gravity), the displacements at any point in the column influence the displacements above it due to load transfer mechanics. Conceptually, this effect continues sequentially upward along the column length. However, in the mathematical formulation of the FEA, the computation involves solving global equations that inherently account for all inter-dependencies between elements and boundary conditions simultaneously, not sequentially. Accordingly, representing columns as directed edges might initially seem appropriate, as it allows separate transmission of force and displacement information. However, in GNNs, undirected graphs are commonly implemented by duplicating directed edges and reversing their directions to ensure bidirectional information flow. Given this practical implementation in GNNs, both beams and columns are treated as undirected edges in our graph model. This approach simplifies the representation while preserving the necessary information flow for structural analysis.
V-A3 Input attributes
In our framework, gravity loads and thermal loads are applied to beams or columns, which dictate the information encapsulated within the nodes and edges of the graph representation.
Node Attributes:
The information associated with each node $i$ in the structure includes geometric details, specifically the 3D coordinates $(x_{i},y_{i},z_{i})$. These coordinates are essential for representing the spatial configuration of the structure. Additionally, since floor height, $h_{i}$, is a critical factor in fire scenarios, it is also included as a key node attribute. To account for fire location, the coordinates of the fire source $(x_{\text{f}},y_{\text{f}},z_{\text{f}})$ and its corresponding floor height $h_{\text{f}}$ are incorporated into the node attributes. To enhance the model’s ability to capture fire-related information, we compute the differences between the node coordinates and its floor height on one hand and the fire location coordinates and its floor height on the other hand, i.e., $\left(x_{i}-x_{\text{f}},y_{i}-y_{\text{f}},z_{i}-z_{\text{f}},h_{i}-h_{\text{f}}\right)$. Additionally, the Euclidean distance between the $i$-th node and the fire location is calculated according to Equation (5) with $x=x_{i}$, $y=y_{i}$, and $z=z_{i}$. These features collectively form the initial attributes of each node $i$, denoted as $\boldsymbol{v}_{i}^{\text{init}}$, comprising a total of 13 features.
Edge Attributes:
  1. 1.
    Material properties: Young’s modulus at ambient temperature, yield strength, and strain-hardening ratio.
  2. 2.
    Geometric properties: Length, floor level, and orientation. Note that the orientation is encoded using a one-hot representation to indicate alignment along the $x$-axis, $y$-axis, or $z$-axis.
  3. 3.
    Gravity loads: Magnitude of applied load on the element.
The initial attributes of each edge $(i,j)$, denoted as $\boldsymbol{e}_{i,j}^{\text{init}}$, are represented by 9-dimensional vector combining these three sets of attributes.
Feature Encoding:
The raw node and edge attributes are not directly input into the GNN. Instead, encoders are employed to map the initial attributes into higher-dimensional spaces, which improve the model’s capacity for feature representation:
  • Node encoder: Extends the 13-dimensional node attributes to $\boldsymbol{v}_{i}^{0}$ with a higher dimensionality, e.g., 64 dimensions. The superscript 00 indicates it is the input of the first GNN layer.
  • Edge encoder: Maps the 9-dimensional edge attributes to $\boldsymbol{e}_{i,j}$ with a higher-dimensional representation, e.g., 32 dimensions.
These encoders, implemented as simple single-hidden-layer MLPs, ensures that the feature dimensions across different layers remain consistent. This consistency is crucial because the number of GNN layers a structure passes through is determined by its maximum number of stories, consistent dimensionality allows for seamless information flow and extraction across layers. This hierarchical encoding approach ensures that the GNN effectively captures the structural and fire-related information necessary for accurate MIDR predictions.
V-A4 Edge update mechanism
Traditional GNNs often do not update edge attributes, $\boldsymbol{e}_{i,j}$, during the iterative update process in each layer. However, in the context of fire scenarios, the attributes of beams and columns (e.g., material properties and geometric characteristics) evolve with temperature changes, directly affecting the computation of the MIDR. To address this limitation, we introduce an EU module that dynamically updates edge attributes during each layer’s computation. After updating node attributes via message passing in each layer of the GNN, the EU module utilizes the attributes of the two adjacent nodes and the edge itself to update the edge attributes. This ensures that the edge features reflect the ongoing structural changes due to fire conditions.
$\boldsymbol{v}_{i}^{k}=\gamma^{k}\left(\boldsymbol{v}_{i}^{k-1},\bigoplus_{j\in\mathcal{N}(i)}\phi^{k}\left(\boldsymbol{v}_{i}^{k-1},\boldsymbol{v}_{j}^{k-1},\boldsymbol{e}_{j,i}^{k-1}\right)\right),$ (9)
The EU module updates the edge attributes as follows:
$\boldsymbol{e}_{j,i}^{k}=\zeta\left(\boldsymbol{v}_{i}^{k},\boldsymbol{v}_{j}^{k}\right)+\boldsymbol{e}_{j,i}^{k-1}.$ (10)
where $\zeta(\cdot)$ is the update function for edge attributes, implemented as a single-hidden-layer MLP.
The inclusion of the EU module enables:
  1. 1.
    Dynamic adaptation: Capture temperature-induced changes in beam and column properties during each layer’s computation.
  2. 2.
    Enhanced modeling: Simultaneously updates node and edge attributes, resulting in a more accurate representation of the structural state.
  3. 3.
    Improved performance: Provides the GNN with the ability to better model the structural transformations under fire scenarios, leading to more precise MIDR predictions.
V-B TL-based Network Training
The proposed GNN framework computes an embedding vector for each node. To predict the IDR, an additional MLP is employed to map the node embedding vector to the corresponding IDR value. We explore two baseline training and inference methods, referred to as Strawman 1 and Strawman 2, as illustrated in Figure 7:
  • Strawman 1: This approach aggregates all node embeddings in a graph into a single graph embedding using a pooling operation. The graph embedding is then used to predict the structure’s MIDR.
  • Strawman 2: In this approach, the IDR for each node is directly predicted using its embedding, and the maximum value is selected as the structure’s MIDR.
Figure 7: Comparison of different training strategies. Head-1 and Head-2 represent the MLPs for different tasks, such as predicting the IDR of each node in a structure and the MIDR of a structure.
While Strawman 1 focuses solely on predicting the MIDR, it fails to leverage the IDR labels available for individual nodes data. Consequently, this approach risks overfitting when the dataset is small. On the other hand, Strawman 2 makes use of all node-level IDR labels, but the inclusion of non-MIDR nodes during training may dilute the model’s focus on predicting the MIDR. To overcome these limitations, we propose a TL-based training method that integrates the strengths of both approaches. By framing the problem as two related tasks –predicting node-level IDRs and structure-level MIDR– we adopt a multi-task TL approach with a two-step training procedure, as illustrated in Figure 7.
In the first step, the GNN is trained to predict the IDR of individual nodes, using an MLP referred to as task head-1, used in Strawman 2, Figure 7. During this step, the GNN learns to extract global structural features. The loss function for this step is defined as follows:
where $d_{m,n,i}^{\,\text{gt}}$ is the ground truth IDR for node $i\in\left(1,2,\dots,I_{m}\right)$ in structure $m\in(1,2,\dots,M)$ computed via OpenSeesRT under fire scenario $n\in\left(1,2,\dots,N_{m}\right)$, $\widehat{d}_{m,n,i}$ represents the predicted IDR from the model, $\Theta_{\text{GNN}}$ is the set of parameters of the GNN, and $\Theta_{\text{head-1}}$ is the set of parameters of task head-1. The optimization objective for this step is as follows:
In the second step, the trained GNN parameters from the first step are reused as a feature extractor. The MLP for node-level IDR prediction (task head-1) is replaced with a new MLP, referred to as task head-2, used in Strawman 1, Figure 7, to predict the MIDR of the structure. During this step, task head-2 is first trained independently, followed by fine-tuning the entire GNN. We define the MIDR for the $m$-th structure under fire scenario $n$ as follows:
$d_{m,n,\max}\triangleq\max_{i=1,2,\dots,I_{m}}d_{m,n,i}.$ (13)
The loss function for this step is defined as follows:
where $d_{m,n,\max}^{\,\text{gt}}$ denotes the ground truth MIDR for the $m$-th structure under fire scenario $n$ computed using OpenSeesRT, $\widehat{d}_{m,n,\max}$ represents the predicted MIDR from the model, and $\Theta_{\text{head-2}}$ is the set of parameters of task head-2. The optimization objective for this step is as follows:
The proposed TL-based methods offers these benefits:
  1. 1.
    Effective utilization of data: Leverage node-level IDR labels while focusing on MIDR prediction.
  2. 2.
    Improved generalization: Enhances the model’s ability to generalize across diverse fire scenarios and structural configurations.
  3. 3.
    Accurate MIDR predictions: Integrates global structural information for precise MIDR predictions, outperforming the individual Strawman methods.
VI Most Fire-Sensitive Point Predictor
With the training of the MIDR Predictor complete, we have developed a neural network-based agent model that effectively serves as a surrogate for the FEA conducted using OpenSeesRT. This agent model presents two critical advantages over traditional FEA: computational efficiency and differentiability. These advantages are leveraged fully in the training of the MFSP Predictor, which is designed to identify the the location within the building structure that exhibits the highest vulnerability under fire conditions, as indicated by the MIDR. Importantly, the training process for the MFSP Predictor no longer depends on direct FEA results. Instead, the fixed parameters of the pretrained MIDR Predictor are utilized as a computational foundation. This allows the MIDR Predictor to remain unchanged during the MFSP Predictor’s training, ensuring consistency and efficiency in the workflow. By relying solely on the outputs of the MIDR Predictor, the MFSP Predictor benefits from the speed and differentiability of the agent model while eliminating the need for repeated computationally expensive FEA simulations.
VI-A Pseudo-Labeling and Loss Choice
As illustrated in Figure 1, the MFSP Predictor operates as the “argmaxer” of the MIDR Predictor. An intuitive choice for the loss function is to minimize the negative output of the MIDR Predictor, denoted as $L_{\text{MIDR}}$. This loss leverages the differentiable nature of the MIDR Predictor as an agent and is expressed as follows:
$L_{\text{MIDR}}=-\frac{1}{\sum_{m}^{M}N_{m}}\sum_{m}^{M}\sum_{n}^{N_{m}}\widehat{d}_{m,n,\max},$ (16)
where $\widehat{d}_{m,n,\max}$ represents the MIDR predicted by the MIDR Predictor for the $n$-th fire scenario in the $m$-th structure.
This approach conceptually resembles the Actor-Critic framework in Reinforcement Learning (RL) [26]: the MFSP Predictor acts as the “Actor”, proposing actions in the form of fire points, while the MIDR Predictor serves as the “Critic”, evaluating the quality of the Actor’s actions. However, similar to challenges in RL, this method can lead to poor training outcome or convergence issues. The MIDR Predictor is a complex function, and optimizing its output through gradient descent often results in local optima or unstable training. Additionally, the fixed nature of MIDR Predictor limits the ability to introduce randomness for balancing exploration and exploitation, a core mechanism in RL.
To address these issues, we propose an alternative approach inspired by the computational efficiency of the MIDR Predictor. Instead of directly using $L_{\text{MIDR}}$, we use the MIDR Predictor to generate pseudo ground truths for the MFSP. This process is depicted in Figure 8. Specifically, at the granularity of rooms, we consider the center point of each room in a building as a potential fire point. The MIDR Predictor is used to estimate the MIDR for each fire point, and the fire point with the highest MIDR selected as the pseudo ground truth MFSP for the building. This pseudo-labeling approach allows efficient labeling of unlabeled data at room-level granularity.
Figure 8: Procedure of pseudo-labeling for the $m$-th building structure.
For a given $m$-th building structure, let the pseudo ground truth MFSP coordinates be $\left(x_{m}^{\,\text{pgt}},y_{m}^{\,\text{pgt}},z_{m}^{\,\text{pgt}}\right)$, with a corresponding MIDR $\widehat{d}_{m}^{\,\text{pgt}}$. Using this pseudo-labeling, we define the Mean Squared Error (MSE) loss as follows:
$\left.+\left(z_{m}^{\,\text{pgt}}-\widehat{z}_{m}\right)^{2}\right),$
where $\widehat{x}_{m}$, $\widehat{y}_{m}$, and $\widehat{z}_{m}$ are the MFSP Predictor’s outputs for the $m$-th structure.
Since the pseudo-labeling process provides ground truth at room-level granularity, it is beneficial to combine $L_{\text{MIDR}}$ and $L_{\text{MSE}}$ into a hybrid loss function via a weighted sum to train the MFSP Predictor. The combined loss $L_{\text{Hybrid}}$ is defined as follows:
$L_{\text{Hybrid}}=w_{1}\,L_{\text{MIDR}}+w_{2}\,L_{\text{MSE}},$ (18)
where $w_{1}$ and $w_{2}$ are weights balancing the contributions of the two loss components. In this study, we fix $w_{2}=1$, so that $w_{1}$ can be interpreted as a measure of the trust or reliance on the pseudo ground truth MFSP generated by the labeling process: the reliance increases as $w_{1}$ decreases. This hybrid approach combines the strengths of leveraging differentiability through $L_{\text{MIDR}}$ and the guidance provided by pseudo ground truth through $L_{\text{MSE}}$, offering a robust method for training the MFSP Predictor.
VI-B TL with GNN in MIDR Predictor
The MFSP Predictor operates as an “argmaxer” for the MIDR Predictor. While it could be designed as a completely new network and trained from scratch (i.e., De novo training), we propose leveraging the pretrained GNN module from the MIDR Predictor through TL. The GNN module in the MIDR Predictor has already captured comprehensive global structural information, making it an foundation for the MFSP Predictor. Specifically, we reuse the GNN module as a feature extractor, replace the MLP task head with a new one tailored for the MFSP task, and fine-tune the entire model to optimize its performance.
VI-B1 Structural information via TL
By reusing the pretrained GNN module, which encodes essential structural and gravity load features, the MFSP Predictor can focus on learning the relationships necessary to identify the MFSP. This approach reduces training time and enhances model performance by building on the already-learned representations of structural information, rather than starting from scratch.
VI-B2 The virtual fire point method
A unique challenge arises due to the way the GNN module processes inputs. In addition to structural and gravity load data, the GNN also requires a fire location input. Common masking techniques, such as zero-value masking (i.e., ignore certain inputs by setting them to zero) used in computer vision or natural language processing, cannot be directly applied here. This is because the coordinates $(0,0,0)$ represent a valid location –specifically, a fire at the bottom corner of the building structure. Thus, zero-value masking is unsuitable for this context.
To overcome this limitation, we introduce a randomized Virtual Fire Point (VFP) method. During training, a random virtual fire location is assigned to each structure for every iteration, independently of prior iterations. This randomized approach forces the network to learn the global structural features necessary for predicting the MFSP, while effectively ignoring the influence of specific fire locations.
VI-B3 Generalization and inference
During inference, the network no longer relies on actual fire locations but uses a predefined VFP to generalize its predictions. For evaluation purposes, we set the VSP to the geometric center of the structure. This ensures that the MSFP Predictor focuses on the overall structural information rather than being biased by the presence of specific fire locations. By combining TL and the VFP method, the MFSP Predictor achieves robust generalization and improved accuracy in identifying the MFSP within diverse structural configurations.
VII Model Implementation and Evaluation
@@ VII-A1 Implementation details
To explore the impact of model capacity on performance, we implemented two versions of the MIDR Predictor: the small predictor and the large predictor. Both models share an identical network architecture but differ in the dimensions of the embedding vectors used for graph nodes and edges. Specifically, the small predictor employs an embedding dimension of $32$, while the large predictor employs an embedding dimension of $64$. These variations allow us to evaluate the trade-offs between model complexity, computational efficiency, and predictive accuracy.
The detailed configurations of each network component, all utilizing ReLU as the activation function in their MLP layers, are outlined below:
  • Initial Encoders for Nodes and Edges: Two distinct single-hidden-layer MLPs serve as encoders for nodes and edges, each tailored to its respective input features. The dimensions of their hidden layers are set to be the same as the output dimension. Consequently, for the small predictor, where both the node and edge embedding dimensions are set to $32$, the node encoder layers are configured as $[13,32,32]$ for the input features, hidden layer, and output layer, respectively. Similarly, the edge encoder layers are $[9,32,32]$. Refer to Section V-A3 for the input attribute counts. For the large predictor, the embedding dimensions are increased to $64$, and the layer configurations are adjusted accordingly. This design ensures the encoders effectively process distinct feature sets for nodes and edges while seamlessly integrating with the GNN architecture.
  • Message Function, $\phi^{k}(\cdot)$: A single-hidden-layer MLP is used for the message function. Its input is the concatenation of the embeddings of the source node and the edge. The output dimension matches the node embedding dimension. For the small predictor, the hidden layer dimension is $64$, while for the large predictor, it is $128$.
@@
  • Update Function, $\gamma^{k}(\cdot)$: The update function combines the previous node embedding with the processed aggregated message. Specifically, the aggregated message for node $i$, denoted as $\boldsymbol{\tilde{v}}_{i}^{k}$, is processed through a single-hidden-layer MLP, $\tilde{\gamma}^{k}(\cdot)$, whose output dimension matches the node embedding dimension. The update rule is given as follows:
@@
    For the small predictor, the hidden layer dimension of $\tilde{\gamma}^{k}(\cdot)$ is $64$, while for the large predictor, it is $128$.
  • Edge Update (EU) Function, $\zeta\left(\cdot\right)$: A single-hidden-layer MLP is used for updating edge attributes. The input is the concatenation of the two adjacent nodes. Thus, the input dimension is twice the node embedding dimension. The output dimension corresponds to the edge embedding dimension. For the small predictor, the hidden layer dimension is $32$, while for the large predictor, it is $64$.
  • Task Head-1, Task Head-2, and Pooling Function: Task heads are implemented as linear layers. The pooling function combines node embedding into a graph embedding by concatenating the mean and max pooling results. Consequently, the input dimension for task head-1 is twice the node embedding dimension, while for task head-2, it equals the node embedding dimension. Recall that task head-1 and task head-2 predict the IDR as a node-level outcome and MIDR as a structure-level outcome, respectively, i.e., a scalar value is output as the final prediction for each head.
The number of layers in the GNN is aligned with the number of stories in the building structure. As the maximum number of the stories of the considered buildings in this study is 7 (refer to Section IV-A), we set the GNN to possess 7 layers. Note that to process a $k$-story structure, only first $k$ layers of the entire GNN is used (refer to Section V-A2). For the small predictor configuration, the total parameter count is approximately $7.1\times 10^{4}$, while for the large predictor configuration, it reaches about $2.8\times 10^{5}$. In the absence of the EU functionality, the parameter counts are marginally reduced to approximately $6.8\times 10^{4}$ for the small predictor and $2.7\times 10^{5}$ for the large predictor.

### v5

@@ VI-F Summary of the Dataset Generation
Figure 8: Histograms of MIDR for labeled and unlabeled building structures with gravity and fire load cases.
This dataset provides the basis for training and testing the performance of the GNN-based framework. Although we employed a simplified rule-based thermal load generation method compared with conventional CFD-based simulations, the temperature field, the changes of the material properties, and the response of the structures, are all still highly nonlinear and complex. Therefore, it is still a challenging task for the NN to predict the MIDRs based on this dataset.
Remark: The dataset generation method involves many simplifications that may not all be representative of the final designs of real-world buildings. However, as stated in Section I, the main usage of the MFSP predictor is in the preliminary design stage where several conceptual designs are explored in an iterative manner and these designs may not be initially in complete compliance with all code provisions. Therefore, although the dataset may contain some unrealistic design cases, it can be viewed as the initial conceptual designs to be evaluated and therefore is valid in evaluating the effectiveness of the intended purpose of the proposed framework.
VII Implementation & Evaluation
@@ VII-A1 Implementation details
To explore the impact of model capacity on performance, we implemente two versions of the MIDR predictor: the small predictor and the large predictor. Both models share an identical network architecture but differ in the dimensions of the embedding vectors used for graph nodes and edges. The architecture and hyper-parameters are chosen with a trial-and-error method and “min-max” normalization is utilized for data-scaling. The small and large predictors employ embedding dimensions of $32$ and $64$, respectively. These variations allow for evaluating trade-offs between model complexity, computational efficiency, and predictive accuracy. The detailed configurations of each network component, all utilizing ReLU as the activation function in their MLP layers, are outlined below:
  • Initial encoders for nodes and edges: Two distinct single-hidden-layer MLPs serve as encoders for nodes and edges, each tailored to its respective input features. The dimensions of their hidden layers are the same as the output dimension. Consequently, for the small predictor, with both node and edge embedding dimensions set to $32$, the node encoder layers are configured as $[13,32,32]$ for the input features, hidden layer, and output layer, respectively. Similarly, the edge encoder layers are $[9,32,32]$. Refer to Section IV-A2 for the input attribute counts. For the large predictor, the embedding dimensions are increased to $64$, and the layer configurations are adjusted accordingly. This design ensures the encoders effectively process distinct feature sets for nodes and edges while seamlessly integrating with the GNN architecture.
  • Message function, $\phi^{k}(\cdot)$: A single-hidden-layer MLP is used for the message function. Its input is the concatenation of the embeddings of the source node and the edge. The output dimension matches the node embedding dimension. The hidden layer dimension for the small and large predictors is $64$ and $128$, respectively.
@@
  • Update function, $\gamma^{k}(\cdot)$: It combines the previous node embedding with the processed aggregated message. Specifically, the aggregated message for node $i$, $\boldsymbol{\tilde{v}}_{i}^{k}$, is processed through a single-hidden-layer MLP, $\tilde{\gamma}^{k}(\cdot)$, whose output dimension matches the node embedding dimension. The update rule is as follows:
@@
    The hidden layer dimension of $\tilde{\gamma}^{k}(\cdot)$ for the small and large predictors is $64$ and $128$, respectively.
  • EU function, $\zeta\left(\cdot\right)$: A single-hidden-layer MLP is used for updating edge attributes. The input is the concatenation of the two adjacent nodes. Thus, the input dimension is twice the node embedding dimension. The output dimension corresponds to the edge embedding dimension. The hidden layer dimension for the small and large predictors is set as $32$ and $64$, respectively.
  • Task head-1, task head-2 & pooling function: Task heads are implemented as linear layers. The pooling function combines node embedding into a graph embedding by concatenating the mean and max pooling results. Consequently, the input dimension for task head-1 (for Strawman 2) is twice the node embedding dimension, while for task head-2 (for Strawman 1), it equals the node embedding dimension. Recall that task head-1 and task head-2 predict the IDR as a node-level outcome and MIDR as a structure-level outcome, respectively, i.e., a scalar value is output as the final prediction for each head.
The number of layers in the GNN is aligned with the number of stories in the building structure. As the maximum number of stories of the considered buildings in this study is 7 (refer to Section VI-A), we set the GNN to possess 7 layers. Note that to process a $k$-story structure, only first $k$ layers of the entire GNN is used (refer to Section IV-A3). For the small predictor configuration, the total parameter count is approximately $7.1\times 10^{4}$, while for the large predictor configuration, it reaches about $2.8\times 10^{5}$. In the absence of the EU functionality, the parameter counts are marginally reduced to approximately $6.8\times 10^{4}$ for the small predictor and $2.7\times 10^{5}$ for the large predictor.